# references.bib

### v1

@@ Abstract
1 Department of Physics, Faculty of Mathematics and Physics,
University of Ljubljana, Jadranska 19, SI-1000 Ljubljana, Slovenia
2 Institute of Mathematics, Physics and Mechanics, Jadranska 19, SI-1000, Ljubljana, Slovenia
chiara.paletta@fmf.uni-lj.si
@@ 1 Introduction
Having an exactly solved model offers a significant advantage, as it often characterizes the universal behaviour of a broader class of potentially unsolvable models, thereby providing the most precise understanding of physical reality. What characterizes the solvability in certain models is the presence of numerous conserved quantities vastly restricting their dynamics and allowing for exact solutions. While this is expected in isolated systems of non-interacting particles, it is remarkable that this phenomenon can also occur in certain interacting theories. The property of having enough – a complete set of – conserved quantities is one of the possible definitions of a quantum integrable model. However, there are various possible precise definitions of integrability [caux2011remarks], and our main focus in this work is on Yang-Baxter integrable models, in particular on Yang-Baxter integrable quantum spin chains. Those are characterized by an $R$-matrix solution of the Yang-Baxter equation (YBE) generating the system’s time evolution.
@@
The $R$-matrix (in appropriate representation) can be used as a building block of the transfer matrix, an object used to construct all the relevant conserved charges characterizing the model. This commutation properties of the transfer matrix, together with the YBE, is the cornerstone of the algebraic Bethe ansatz [faddeev1996algebraic], a method for diagonalising integrable many-body Hamiltonians. This exact technique provides also insightful information on the calculation of correlation functions. When the spin chain has open boundary conditions, to define integrability, together with the standard YBE, two extra relations need to be considered, [Sklyanin:1988yz, NepomechieJPA:1991, nepomechie2018surveying]. Those describe, respectively, the scattering of particles with the right and left boundaries. The solutions of these relations are referred to as reflection matrices (or $K$-matrices). The technique of algebraic Bethe ansatz generally extends for this class of model, see e.g. Ref. [nepomechie2003bethe].
@@
A many body operator $M$ will be referred to as a quantum circuit when it can be factorized into a finite sequence of 2-particle operators, the so-called gates. An integrable quantum circuit may be realized when $M$ commutes with a transfer matrix of an integrable model. Based on this idea, the integrable trotterization procedure was developed. The idea originates from a work by Baxter [baxter2016exactly], where it was understood that the values of the so-called inhomogeneities of the transfer matrix of the six-vertex model can be fixed to define a discrete time parallel update on a periodic lattice. This idea has been elaborated in [vanicat2018integrable, vanicat2018integrable2], by providing a general framework (independent of the $R$-matrix considered) to define an integrable unitary discrete time evolution. The trotterization procedure works with either periodic and open boundary conditions and in both cases it appears to be two-step Floquet (time-periodic) dynamics.
The advantage of having an integrable quantum circuit relies on the possibility to use the integrability technique to compute the spectrum. Those analytical results may be used for the calibration and error mitigations in modern engineered quantum platforms, [circuitBA1, circuitBA2]. The setup of integrable quantum circuits has also been instrumental for demonstration of universal superdiffusive scaling aspects of quantum dynamics, [circuitBA3]. Recently, in the setup of integrable quantum circuits, correlation functions of strings of spin operators were computed, [circuitBA4].
@@
An interesting class of quantum circuits that we discuss in this work, is the one where the quantum gates operate on density operators directly rather than on the pure quantum states. Using the vectorization (aka thermofield double) representation, we can consider gates which act on the tensor products of pairs of local Hilbert spaces. A general Krauss representation of time evolution of a density matrix then results in a non-unitary quantum circuit. In Ref. [sa2021integrable] it was demonstrated that trotterization of the Hubbard model with imaginary interaction strength corresponds to an integrable open quantum circuit with a dephasing noise. This is an example of the general mapping between Liouvillians of open many-body systems and Bethe-ansatz integrable systems on (thermofield) doubled Hilbert spaces [medvedyeva2016exact, ziolkowska2020yang, classificationlind, de2022bethe].
The quantum gate that we consider in this work is given by the tensor product of two elementary quantum gates. This corresponds, in the continuum time evolution, to the coherent evolution part of the Lindblad master equation, [breuer2002theory, manzano2020short].
@@
$\dot{\rho}(t)=\frac{d}{dt}\rho=\mathcal{L}\rho:=i[\rho,H]+\Gamma\sum_{j}\Big{[}\ell_{j}\rho\ell_{j}^{\dagger}-\frac{1}{2}\{\ell_{j}^{\dagger}\ell_{j},\rho\}\Big{]},$ (1.1)
@@
The stationary state of this model is the so-called non equilibrium steady state (NESS) and satisfies $\mathcal{L}\rho_{\rm NESS}=0$. In 2011, the NESS density operator of the boundary driven Lindblad master equation for the XXZ spin 1/2 chain has been solved in terms of matrix product operator ansatz for particular spin source/sink boundary operators, [prosen2011open, prosen2011exact]. For a review see [prosen2015matrix]. Initially, these solutions appeared unrelated to the conventional theory of quantum integrability. However, later it was understood that they are connected to the infinite-dimensional solutions of YBE associated with non-unitary irreducible representations of the model’s quantum group symmetry, [ilievski2014exact, ilievski2014quantum, prosen-exterior-lindblad]. Specifically, a link was found between the matrix product form of the non-equilibrium steady states and the integrable structure of the bulk Hamiltonian. It is also possible to write exact solutions for the NESS of a quantum XX chain in the presence of bulk dephasing and arbitrary local boundary spin driving [vznidarivc2010exact]. Another class of models for which, moreover, the full spectrum and eigenvectors were calculated is the XY spin chain with jump operator being linear in canonical fermionic operators, [prosen2008third].
@@ 1.1 Short summary of the paper
This paper is organized as follows. In section 2, we review key concepts of quantum integrability for spin chains with both periodic [faddeev1996algebraic] and open boundary conditions. We detail the Sklyanin construction [Sklyanin:1988yz] for the left and right reflection algebras and summarize the protocol for constructing integrable quantum circuits with two-step Floquet dynamics, [vanicat2018integrable, vanicat2018integrable2]. From section 3, the original part of the paper starts. Initially, we define the two-replica quantum circuits coupled through the boundary and explain the method for solving the boundary Yang-Baxter equations. While we use general open boundary conditions, the quantum elementary gate of the bulk is given by the tensor product of pairs two qubit gates. A central question we investigate is whether the quantum gates corresponding to the left and right boundaries exhibit a non-factorized expression. In section 4, we report our results. We begin by examining several integrable models: interacting models such as the Heisenberg XXX and XXZ spin chains, as well as the free fermion XX spin chain. For each model, we discuss the solution of the Sklyanin reflection algebra in the doubled (two replica) Hilbert space. Our main finding is that the solution of the reflection algebra is non-factorized only in the case where the bulk is the free fermion XX spin chain. In that case, the boundary interaction can be non-trivial and the model may represent an interesting example of integrable unitary quantum dynamics in its own right. In section 5, we then discuss the implications of our results for continuous time and open system dynamics. The conserved charges of the models are derived from the higher-order derivatives of the transfer matrix. We analyze under which conditions the boundary terms can be expressed as a Lindbladian evolution. We conclude that this identification is only possible if the operators at the left and right boundaries of the spin chain act as sources of particle injection or removal. This implies that the Yang-Baxter integrability property of a model (both in the bulk and in the boundaries) imposes stricter conditions than those required for finding the analytical expression of the NESS of the model. While finding the NESS is possible for a broader class of models [prosen2011exact], the Yang-Baxter integrability property is more restrictive.
@@ 2.1.1 Periodic Boundary Condition
A quantum integrable model has an infinite number of conserved charges and it is characterized by an $R$-matrix solution of the Yang-Baxter equation (YBE), [baxter2016exactly]
$R_{12}(u)R_{13}(u+v)R_{23}(v)=R_{23}(v)R_{13}(u+v)R_{12}(u).$ (2.1)
This is a matrix relation defined in the triple tensor product vector space $\text{End}(V\otimes V\otimes V)$, with $V\equiv\mathbb{C}^{n}$ the $n$-th dimensional complex vector space. The $R$-matrix is defined in $\text{End}(V\otimes V)$ and in the indexed version $R_{ij}$, the subscripts denote which of the three spaces $R$ acts on nontrivially (for example $R_{12}=R\otimes\mathbb{I}$, $\mathbb{I}$ being the identity operator in $\mathbb{C}^{n}$) and $u,v$ are known as spectral parameters and can take values in $\mathbb{C}$. At both classical [^1] The classical YBE takes a different expression than (2.1), we refer to [torrielli2016classical, analecture] for a detailed explanation. and quantum levels, the YBE is considered the crucial component of integrability.
@@
$t(u,\mathbf{w})=\text{tr}_{0}\,T_{0}(u,\mathbf{w}),$ (2.4)
@@
By using the RTT relation (2.2), it follows that
$[t(u,\mathbf{w}),t(v,\mathbf{w})]=0.$ (2.5)
$t(u,\mathbf{w})$ can be considered the generating function of the conserved charges $\mathbb{Q}_{j}$ of a quantum system. In fact, expanding the logarithm of $t(u,\mathbf{w})$ in a power series and calling $\mathbb{Q}_{i}$ the coefficient of $u^{i}$,
$\mathbb{Q}_{i}=\partial^{i-1}_{u}\log t(u,\mathbf{w})|_{u=0}$ (2.6)
@@
$[\mathbb{Q}_{i},\mathbb{Q}_{j}]=0,\quad i,j=1,2\ldots.$ (2.7)
The involutivity condition holds for any values of the inhomogeneities $\mathbf{w}$. In many practical cases, $\mathbf{w}=(0,0,\dots,0)$ and all charges $\mathbb{Q}_{j}$ are translationally invariant sums of local operators of interaction range $j$, if the R-matrix obey the regularity condition $R(0)=P$ where $P$ is a permutation (see below). For values of $w_{i}$ close to 00, the charges are typically quasi-local, [ilievski2016quasilocal].
Usually, when considering a physical model, we are interested in its dynamics. For the continuous time evolution, the operator that generates the dynamics is the Hamiltonian $\mathbb{Q}_{2}=H$, connected to the transfer matrix via [^2] As will be clarified, this condition is $H=\partial_{u}t(u,\mathbf{0})$ for a spin chain with open boundary condition.
$H=\partial_{u}\log t(u,\mathbf{0})|_{u=0}.$ (2.8)
In this work, we first focus in the discrete time evolution. The $R$-matrix of an integrable model can be used as the quantum gate to build an integrable quantum circuit. In section 2.2, we review the trotterization procedure [vanicat2018integrable, vanicat2018integrable2]: the inhomogeneities can be fixed to define a two-steps Floquet dynamics that preserves integrability for both periodic and open boundary conditions.
@@ 2.1.2 Open Boundary Condition
For an integrable model with open boundary conditions, other than the usual Yang-Baxter equation (2.1), we have to introduce two new matrices $K^{R}(u)$ and $K^{L}(u)$ that take into account the right and left boundary of the spin chain. They are called reflection matrices or, in the same spirit of Eqs. (2.2), reflection algebra.
To avoid confusion, we indicate the $K$-matrices as $K_{a}^{b}$, where $a$ indicates the site where the matrix is acting non-trivially, while the label $b\in\{L,R\}$ denotes the left ($L$) or right ($R$) boundary matrix. Extra superscript may indicate the inverse of the matrix ($-1$) or the transpose ($t$).
@@
Figure 1: Graphical representation of the right reflection algebra (2.9).
@@
$R_{12}(-u+v){{K}_{1}^{L}}^{t_{1}}(u)R_{21}(-u-v-2\eta){{K}_{2}^{L}}^{t_{2}}(v)={{K}_{2}^{L}}^{t_{2}}(v)R_{21}(-u-v-2\eta){{K}_{1}^{L}}^{t_{1}}(u)R_{12}(-u+v).$ (2.10)
The solutions of (2.9) and (2.10) are related by one of the following automorphisms
@@
It can be shown that the three automorphism are invertible and that by plugging the mapping for $K^{R}/K^{L}$ into (2.9)/(2.10), one obtains (2.10)/(2.9). In this way, we can conclude that all the solutions of $K^{L}$ are related to the ones for $K^{R}$.
By following Sklyanin’s construction [Sklyanin:1988yz], the generator of the conserved charges is the double row transfer matrix
$t(u)=\text{tr}_{0}\Big{(}K_{0}^{L}(u)T(u)K_{0}^{R}(u)\hat{T}(u)\Big{)},$ (2.11)
@@
Graphically, we can think about the process as depicted in Fig. 2, [bajnok2006equivalences].
Figure 2: Double row transfer matrix (2.11).
This describes an auxiliary particle scattering through all the others, hitting the left wall, reflecting with opposite momenta, and after scattering all the others again, hitting the right wall [^4] We remark that, to simplify the notation, we are referring to $K^{L}(u)$ as the solution of the dual reflection equation (2.10). This is the object entering in the Fig. 2. If, on the other hand, we were to draw the equivalent of Fig. 1 for the left boundary, we would obtain an equation for the left reflection algebra ($\tilde{K}^{L}(u)$). The two matrices are related by a transformation as shown in [vanicat2018integrable], specifically, ${K}^{L}_{1}(u)=\text{tr}_{0}\Big{(}\tilde{K}^{L}_{0}(-u)R_{01}(-2u)P_{01}\Big{)}$..
By using the Yang-Baxter equation (2.1) and the reflection equations (2.9) and (2.10), one can prove that
$[t(u),t(v)]=0.$ (2.14)
@@
$t(u,\mathbf{w})\,=\,$ $\text{tr}_{0}\Big{(}K_{0}^{L}(u)T(u,\mathbf{w})K_{0}^{R}(u)\hat{T}(u,\mathbf{w})\Big{)},$ (2.15)
@@
$T(u,\mathbf{w})=R_{0L}(u-w_{L})R_{0,L-1}(u-w_{L-1})\cdots R_{01}(u-w_{1}),$ (2.16)
$\hat{T}(u,\mathbf{w})=R_{10}(u+w_{1})R_{20}(u+w_{2})\cdots R_{L0}(u+w_{L}).$ (2.17)
@@
$[t(u,\mathbf{w}),t(v,\mathbf{w})]=0\,.$ (2.18)
As for periodic boundary condition, for the continuous time dynamics, the derivative $\partial_{u}t(u,\mathbf{0})|_{u=0}$ defines a quantum spin chain Hamiltonian. In the next section, strictly following [vanicat2018integrable], we show how to fix the values of the inhomogeneities to build a discrete time process.
We remark that, the boundary equation for $K^{L}$ (2.10) holds if the $R$-matrix satisfies the crossing unitarity property $R^{t_{1}}(u)R^{t_{1}}(-u-2\eta)\propto\mathbb{I}$. This can be understood by following the proof by Sklyanin, in [Sklyanin:1988yz].
In the absence of this property, one can follow the same proof and only use the unitarity property. In this way, the dual reflection equation for $K^{L}$ is, [vanicat2018integrable]
$R_{12}(-u+v){{K}_{1}^{L}}^{t_{1}}(u)\Big{(}\big{(}R_{21}^{t_{2}}\big{)}^{-1}\Big{)}^{t_{2}}(u+v){{K}_{2}^{L}}^{t_{2}}(v)={{K}_{2}^{L}}^{t_{2}}(v)\Big{(}\big{(}R_{21}^{t_{2}}\big{)}^{-1}\Big{)}^{t_{2}}(u+v){{K}_{1}^{L}}^{t_{1}}(u)R_{12}(-u+v).$ (2.19)

### v2

# Integrability of open boundary driven quantum circuits

@@ Abstract
Copyright attribution to authors.
This work is a submission to SciPost Physics.
License information to appear upon publication.
Publication information to appear upon publication.
Received Date
Accepted Date
Published Date
@@ 1 Introduction
Having an exactly solved model offers a significant advantage, as it often characterizes the universal behaviour of a broader class of potentially unsolvable models, thereby providing the most precise understanding of physical reality. What characterizes the solvability in certain models is the presence of numerous conserved quantities vastly restricting their dynamics and allowing for exact solutions. While this is expected in isolated systems of non-interacting particles, it is remarkable that this phenomenon can also occur in certain interacting theories. The property of having enough – a complete set of – conserved quantities is one of the possible definitions of a quantum integrable model. However, there are various possible precise definitions of integrability [1], and our main focus in this work is on Yang-Baxter integrable models, in particular on Yang-Baxter integrable quantum spin chains. Those are characterized by an $R$-matrix solution of the Yang-Baxter equation (YBE) generating the system’s time evolution.
@@
The $R$-matrix (in appropriate representation) can be used as a building block of the transfer matrix, an object used to construct all the relevant conserved charges characterizing the model. This commutation properties of the transfer matrix, together with the YBE, is the cornerstone of the algebraic Bethe ansatz [2], a method for diagonalising integrable many-body Hamiltonians. This exact technique provides also insightful information on the calculation of correlation functions. When the spin chain has open boundary conditions, to define integrability, together with the standard YBE, two extra relations need to be considered, [3, 4, 5]. Those describe, respectively, the scattering of particles with the right and left boundaries. The solutions of these relations are referred to as reflection matrices (or $K$-matrices). The technique of algebraic Bethe ansatz generally extends for this class of model, see e.g. Ref. [6].
@@
A many body operator $M$ will be referred to as a quantum circuit when it can be factorized into a finite sequence of 2-particle operators, the so-called gates. An integrable quantum circuit may be realized when $M$ commutes with a transfer matrix of an integrable model. Based on this idea, the integrable trotterization procedure was developed. The idea originates from a work by Baxter [7], where it was understood that the values of the so-called inhomogeneities of the transfer matrix of the six-vertex model can be fixed to define a discrete time parallel update on a periodic lattice. This idea has been elaborated in [8, 9], by providing a general framework (independent of the $R$-matrix considered) to define an integrable unitary discrete time evolution. The trotterization procedure works with either periodic and open boundary conditions and in both cases it appears to be two-step Floquet (time-periodic) dynamics.
The advantage of having an integrable quantum circuit relies on the possibility to use the integrability technique to compute the spectrum. Those analytical results may be used for the calibration and error mitigations in modern engineered quantum platforms, [10, 11]. The setup of integrable quantum circuits has also been instrumental for demonstration of universal superdiffusive scaling aspects of quantum dynamics, [12]. Furthermore, in a particular limit, the connection between the Floquet construction and a non-rational conformal field theory was established, [13]. It was also shown that for the quantum circuit constructed from the trotterization of the XXZ spin chain, strong zero modes can be constructed in certain regions of parameter space, [14]. Recently, in the setup of integrable quantum circuits, correlation functions of strings of spin operators were computed, [15].
@@
An interesting class of quantum circuits that we discuss in this work, is the one where the quantum gates operate on density operators directly rather than on the pure quantum states. Using the vectorization (aka thermofield double) representation, we can consider gates which act on the tensor products of pairs of local Hilbert spaces. A general Krauss representation of time evolution of a density matrix then results in a non-unitary quantum circuit. In Ref. [16] it was demonstrated that trotterization of the Hubbard model with imaginary interaction strength corresponds to an integrable open quantum circuit with a dephasing noise. This is an example of the general mapping between Liouvillians of open many-body systems and Bethe-ansatz integrable systems on (thermofield) doubled Hilbert spaces [17, 18, 19, 20].
The quantum gate that we consider in this work is given by the tensor product of two elementary quantum gates. This corresponds, in the continuum time evolution, to the coherent evolution part of the Lindblad master equation, [21, 22].
@@
$\dot{\rho}(t)=\frac{d}{dt}\rho=\mathcal{L}\rho:=i[\rho,H]+\Gamma\sum_{j}\Big{[}\ell_{j}\rho\ell_{j}^{\dagger}-\frac{1}{2}\{\ell_{j}^{\dagger}\ell_{j},\rho\}\Big{]},$ (1)
@@
The stationary state of this model is the so-called non equilibrium steady state (NESS) and satisfies $\mathcal{L}\rho_{\rm NESS}=0$. In 2011, the NESS density operator of the boundary driven Lindblad master equation for the XXZ spin 1/2 chain has been solved in terms of matrix product operator ansatz for particular spin source/sink boundary operators, [23, 24]. For a review see [25]. Initially, these solutions appeared unrelated to the conventional theory of quantum integrability. However, later it was understood that they are connected to the infinite-dimensional solutions of YBE associated with non-unitary irreducible representations of the model’s quantum group symmetry, [26, 27, 28]. Specifically, a link was found between the matrix product form of the non-equilibrium steady states and the integrable structure of the bulk Hamiltonian. For a quantum XX chain in the presence of bulk dephasing and arbitrary local boundary spin driving, it is possible to write the solutions for the NESS to the 2nd order in the driving strength of [29]. For the XX spin chain of finite length coupled to reservoirs at both ends, the NESS can be written as an MPO with fixed bound dimension 4 independent of the chain length, [30]. Another class of models for which, moreover, the full spectrum and eigenvectors were calculated is the XY spin chain with jump operator being linear in canonical fermionic operators, [31].
@@ 1.1 Short summary of the paper
This paper is organized as follows. In section 2, we review key concepts of quantum integrability for spin chains with both periodic [2] and open boundary conditions. For the case with open boundary condition, we detail the Sklyanin construction [3] for the left and right reflection algebras and summarize the protocol for constructing integrable quantum circuits with two-step Floquet dynamics, [8, 9]. From section 3, the original part of the paper starts. Initially, we define the two-replica quantum circuits coupled through the boundary and explain the method for solving the boundary Yang-Baxter equations. The quantum elementary gate of the bulk is given by the tensor product of pairs two qubit gates. A central question we investigate is whether the quantum gates corresponding to the left and right boundaries exhibit a non-factorized expression. In section 4, we report our results. We begin by examining several integrable models: interacting models such as the Heisenberg XXX and XXZ spin chains, as well as the free fermion XX spin chain. For each model, we discuss the solution of the Sklyanin reflection algebra in the doubled (two replica) Hilbert space. Our main finding is that the solution of the reflection algebra is non-factorized only in the case where the bulk is the free fermion XX spin chain. In that case, the boundary interaction can be non-trivial and the model, restricting the possible values of the free parameters, may represent an interesting example of integrable unitary quantum dynamics in its own right. In section 5, we then discuss the implications of our results for continuous time and open system dynamics. The conserved charges of the models are derived from the higher-order derivatives of the transfer matrix. We analyze under which conditions the boundary terms can be expressed as a Lindbladian evolution. We conclude that this identification is only possible if the operators at the left and right boundaries of the spin chain act as sources of particle injection or removal. This implies that the Yang-Baxter integrability property of a model (both in the bulk and in the boundaries) imposes stricter conditions than those required for finding the analytical expression of the NESS of the model. While finding the NESS is possible for a broader class of models [24], the Yang-Baxter integrability property is more restrictive.
@@ 2.1.1 Periodic Boundary Condition
A quantum integrable model has an infinite number of conserved charges and it is characterized by an $R$-matrix solution of the Yang-Baxter equation (YBE), [7]
$R_{12}(u)R_{13}(u+v)R_{23}(v)=R_{23}(v)R_{13}(u+v)R_{12}(u).$ (2)
This is a matrix relation defined in the triple tensor product vector space $\text{End}(V\otimes V\otimes V)$, with $V\equiv\mathbb{C}^{n}$ the $n$-th dimensional complex vector space. The $R$-matrix is defined in $\text{End}(V\otimes V)$ and in the indexed version $R_{ij}$, the subscripts denote which of the three spaces $R$ acts on nontrivially (for example $R_{12}=R\otimes\mathbb{I}$, $\mathbb{I}$ being the identity operator in $\mathbb{C}^{n}$) and $u,v$ are known as spectral parameters and can take values in $\mathbb{C}$. At both classical [^1] The classical YBE takes a different expression than (2), we refer to [32, 33] for a detailed explanation. and quantum levels, the YBE is considered the crucial component of integrability.
@@
$t(u,\mathbf{w})=\text{tr}_{0}\,T_{0}(u,\mathbf{w}),$ (5)
@@
By using the RTT relation (3), it follows that
$[t(u,\mathbf{w}),t(v,\mathbf{w})]=0.$ (6)
$t(u,\mathbf{w})$ can be considered the generating function of the conserved charges $\mathbb{Q}_{j}$ of a quantum system. In fact, expanding the logarithm of $t(u,\mathbf{w})$ in a power series and calling $\mathbb{Q}_{i}$ the coefficient of $u^{i{-1}}$,
$\mathbb{Q}_{i}=\partial^{i-1}_{u}\log t(u,\mathbf{w})|_{u=0}$ (7)
@@
$[\mathbb{Q}_{i},\mathbb{Q}_{j}]=0,\quad i,j=1,2\ldots.$ (8)
The involutivity condition holds for any values of the inhomogeneities $\mathbf{w}$. In many practical cases, $\mathbf{w}=\mathbf{0}=(0,0,\dots,0)$ and all charges $\mathbb{Q}_{i}$ are translationally invariant sums of local operators of interaction range $i$, if the R-matrix obey the regularity condition $R(0)=P$ where $P$ is a permutation (see below). For values of $w_{i}$ close to 00, the charges are typically quasi-local, [34].
Usually, when considering a physical model, we are interested in its dynamics. For the continuous time evolution, the operator that generates the dynamics is the Hamiltonian $\mathbb{Q}_{2}=H$, connected to the transfer matrix via [^2] The $\log$ guarantees the locality of the charges for periodic boundary condition. For open boundary condition, (9) is $H=\partial_{u}t(u,\mathbf{0})|{{}_{u=0}}$ for a spin chain with open boundary condition.
$H=\partial_{u}\log t(u,\mathbf{0})|_{u=0}.$ (9)
In this work, we first focus in the discrete time evolution. The $R$-matrix of an integrable model can be used as the quantum gate to build an integrable quantum circuit. In section 2.2, we review the trotterization procedure [8, 9]: the inhomogeneities can be fixed to define a two-steps Floquet dynamics that preserves integrability for both periodic and open boundary conditions.
@@ 2.1.2 Open Boundary Condition
For an integrable model with open boundary conditions, other than the usual Yang-Baxter equation (2), we have to introduce two new matrices $K^{R}(u)$ and $K^{L}(u)$ that take into account the right and left boundary of the spin chain. They are called reflection matrices or, in the same spirit of Eqs. (3), reflection algebra.
To avoid confusion, we indicate the $K$-matrices as $K_{a}^{b}$, where $a$ indicates the site where the matrix is acting non-trivially, while the label $b\in\{L,R\}$ denotes the left ($L$) or right ($R$) boundary matrix. Extra superscript may indicate the inverse of the matrix ($-1$) or the transpose ($t$) or the partial transpose ($t_{1}$) and ($t_{2}$) with respect to the first or to the second space.
@@
Figure 1: Graphical representation of the right reflection algebra (10).
@@
$R_{12}(-u+v){{K}_{1}^{L}}^{t_{1}}(u)R_{21}(-u-v-2\eta){{K}_{2}^{L}}^{t_{2}}(v)={{K}_{2}^{L}}^{t_{2}}(v)R_{21}(-u-v-2\eta){{K}_{1}^{L}}^{t_{1}}(u)R_{12}(-u+v).$ (11)
The solutions of (10) and (11) are related by one of the following automorphisms
@@
It can be shown that the three automorphism are invertible and that by plugging the mapping for $K^{R}/K^{L}$ into (10)/(11), one obtains (11)/(10). In this way, we can conclude that all the solutions of $K^{L}$ are related to the ones for $K^{R}$.
By following Sklyanin’s construction [3], the generator of the conserved charges is the double row transfer matrix
$t(u)=\text{tr}_{0}\Big{(}K_{0}^{L}(u)T(u)K_{0}^{R}(u)\hat{T}(u)\Big{)},$ (12)
@@
Graphically, we can think about the process as depicted in Fig. 2, [35].
Figure 2: Double row transfer matrix (12).
This describes an auxiliary particle scattering through all the others, hitting the left wall, reflecting with opposite momenta, and after scattering all the others again, hitting the right wall [^4] We remark that, to simplify the notation, we are referring to $K^{L}(u)$ as the solution of the dual reflection equation (11). This is the object entering in the Fig. 2. If, on the other hand, we were to draw the equivalent of Fig. 1 for the left boundary, we would obtain an equation for the left reflection algebra ($\tilde{K}^{L}(u)$). The two matrices are related by a transformation as shown in [8], specifically, ${K}^{L}_{1}(u)=\text{tr}_{0}\Big{(}\tilde{K}^{L}_{0}(-u)R_{01}(-2u)P_{01}\Big{)}$..
By using the Yang-Baxter equation (2) and the reflection equations (10) and (11), one can prove that, [3, 36]
$[t(u),t(v)]=0.$ (15)
@@
$t(u,\mathbf{w})\,=\,$ $\text{tr}_{0}\Big{(}K_{0}^{L}(u)T(u,\mathbf{w})K_{0}^{R}(u)\hat{T}(u,\mathbf{w})\Big{)},$ (16)
@@
$T(u,\mathbf{w})=R_{0L}(u-w_{L})R_{0,L-1}(u-w_{L-1})\cdots R_{01}(u-w_{1}),$ (17)
$\hat{T}(u,\mathbf{w})=R_{10}(u+w_{1})R_{20}(u+w_{2})\cdots R_{L0}(u+w_{L}).$ (18)
@@
$[t(u,\mathbf{w}),t(v,\mathbf{w})]=0\,.$ (19)
As for periodic boundary condition, for the continuous time dynamics, the derivative $\partial_{u}t(u,\mathbf{0})|_{u=0}$ defines a quantum spin chain Hamiltonian. In the next section, strictly following [8], we show how to fix the values of the inhomogeneities to build a discrete time process.
We remark that, the boundary equation for $K^{L}$ (11) holds if the $R$-matrix satisfies the crossing unitarity property $R^{t_{1}}(u)R^{t_{1}}(-u-2\eta)\propto\mathbb{I}$. This can be understood by following the proof by Sklyanin, in [3].
In the absence of this property, one can follow the same proof and only use the unitarity property. In this way, the dual reflection equation for $K^{L}$ is [8, 37]
$R_{12}(-u+v){{K}_{1}^{L}}^{t_{1}}(u)\Big{(}\big{(}R_{21}^{t_{2}}\big{)}^{-1}\Big{)}^{t_{2}}(u+v){{K}_{2}^{L}}^{t_{2}}(v)={{K}_{2}^{L}}^{t_{2}}(v)\Big{(}\big{(}R_{21}^{t_{2}}\big{)}^{-1}\Big{)}^{t_{2}}(u+v){{K}_{1}^{L}}^{t_{1}}(u)R_{12}(-u+v).$ (20)